OpenStudy (anonymous):

Find the quotient of 21x^2y^6 + 6xy^3 - 30xy/3xy. 21x2y6 + 6xy3 − 10 7xy5 + 2y2 + 10 7xy5 + 2y2 − 10 7xy5 + 2y − 10

OpenStudy (anonymous):

@mivanov

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Divide each term by 3xy

OpenStudy (anonymous):

the second terms of all answers are different so just look at them don't bother with the rest and do divide them by 3xy

OpenStudy (mertsj):

\[\frac{21x^2y^6}{3xy}+\frac{6xy^3}{3xy}=\]

OpenStudy (anonymous):

hm the whole thing is divided? i mistoke it because i saw no parenthesis

OpenStudy (anonymous):

there is no parenthesis

OpenStudy (mertsj):

Based on the answer choices I think the problem is: \[\frac{21x^2y^6+6xy^3-30xy}{3xy}\]

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

so just divide by 3xy yes

OpenStudy (anonymous):

simple as that

OpenStudy (anonymous):

it is exactly like that

OpenStudy (mertsj):

Which is equal to: \[7xy^5+2y^2-10\]